# Learning a Stable, Safe, Distributed Feedback Controller for a Heterogeneous Platoon of Vehicles

### v1

@@ IV-A Controller learning
IV-B Hardware experiments
To validate our algorithm on hardware, we use the F1Tenth platform, modified as described in [15]. We test the vehicles in a roughly 4 m by 8 m oval racetrack, shown in Fig. 1. We model each vehicle’s dynamics using Eq. 1 with $\tau_{i}=0.3$ (found using a least-squares fit to experimental data). The lead vehicle uses the quadratic cost function DMPC controller described in [15] to track a safe velocity based on the current curvature of the course (because we evaluate against the DMPC method and we want the lead vehicle to act similarly for all experiments). This leads to the lead vehicle assuming faster trajectories during the straight portions of the course and slower trajectories around the curves. Each following vehicle uses the neural network controller learned using algorithm 1 with a desired distance of 0.75 m.
To ensure safety of the learned controller, we use $L_{\text{safe}}$ to penalize when a vehicle has a distance of less than 0.25 m to its predecessor. To encourage smooth trajectories with $L_{\text{slew}}$, we penalize large actions and large changes in speed from one timestep to the next. We also add two ReLU layers to the neural network to ensure the controller only outputs commands such that $\left|u_{i}\right|\leq 3\,\text{ m/s}^{2}$. To guide the controller toward closed-loop stability, we use ${\cal L}_{\text{stab}}$ to penalize a running sum of the difference in error from the next timestep to the current timestep over the horizon (because we ideally want the errors to decrease). Other choices for $L_{\text{stab}}$ can try to guide the controller towards improving the convergence rate of the system to the origin. We ran algorithm 1 with $N=5$, where both $V$ and $\pi$ are two-layer feedforward neural networks with 8 neurons in each hidden layer and leaky ReLU activation functions with negative slope 0.1 for around 100 episodes to learn a controller.
@@ IV-B Hardware experiments
Based on the results in Figs. 3 and 4, it is difficult to determine which algorithm performs the best. We do note that when using the neural network controller, zero collisions occurred between vehicles or with the course boundaries. With the other two controllers, however, vehicles would sometimes collide with the course boundaries. Further, with the linear feedback controller, vehicles would sometimes collide with one another when braking at the end of the experiment. We speculate that the collisions with the boundaries occurred because the controllers attempted to follow the predecessor’s velocity profile too aggressively when rounding the curves, leading to lateral slipping.
@@
IV-C Simulation experiments
@@
For a platoon using a PF topology and a CDH spacing policy, ensuring safety becomes more difficult as the platoon size increases [7, 8, 9]. To investigate the limits of our approach, we simulate a platoon of $N=100$ vehicles and train a controller using a desired distance of 5 m. We use an almost identical setup as in the hardware section, except we use two hidden layers with 32 neurons each, we penalize a vehicle for getting within 2 m of its predecessor, and we use a heterogeneous platoon where the dynamics parameter $\tau_{i}\in[0.2,0.8]$ is selected at random. As mentioned in the previous section, we cannot verify the Lyapunov certificates because 1) the search space becomes very large and 2) the number of binary variables required to solve Eq. 7b increases linearly with the platoon size leading to potentially exponential increase in the time required to solve the MILPs. However, we still find that training for a very small amount of time (around 100 simulated episodes) yields controllers that perform well in test scenarios.
@@
As shown in [15], there is a large gap in the performance of DMPC controllers and linear feedback controllers as platoon size scales. Though DMPC should always outperform feedback methods due to the advantage of being able to communicate planned trajectories, it would be useful in practice if a feedback controller can perform similarly well. For example, in the scenario where communication is temporarily lost, it would be beneficial to have a feedback controller that can take the place of the DMPC method without disrupting the platoon’s motion. Thus, we again compare our learned controller to DMPC and linear feedback, with the goal of bridging the gap between DMPC methods and linear feedback methods.
For the simulation experiments, we have the first vehicle track a velocity profile that starts at 20 m/s, accelerates to 25 m/s, then decelerates back to 20 m/s. We add Gaussian noise to both the forward dynamics and to the sensing of the position/velocity errors to ensure the learned controller still works in a noisy setting. Note that for a vehicle traveling 20 m/s, common practice would suggest a spacing of 40 m [11], so a 5 m spacing is aggressive for this scenario. The simulation results for the neural network controller are shown in Fig. 5. As expected when using a feedback controller with a PF topology, the errors do propagate down the platoon, leading to the vehicles further away needing to overcompensate by speeding up and slowing down more than vehicles earlier in the platoon. However, the controller does stabilize the platoon to the true desired velocities quickly and without any collisions within the platoon.
@@
Figure 5: Simulated trajectories for a platoon of 100 vehicles. The lead vehicle tracks the velocity profile shown by the reference velocity (black line) in the velocity plot. The faded lines in the position plot show the true desired position of each vehicle.
@@
IV-D Discussion
The algorithm proposed in this work and in other prior works do not work well with regard to learning a Lyapunov function for very high-dimensional systems. However, these algorithms used to concurrently learn a Lyapunov function and a controller do generate controllers that can perform better than hand-designed controllers and translate well from simulation to the real world.
To try to improve the convergence rate of our algorithm, we attempted to “warm-start” the Lyapunov function by sharing weights from a previously verified Lyapunov function with smaller $N$. However, we did not find that this improved performance. Instead, convergence rates were very inconsistent even for small $N$. For $N=2$, the algorithm would often converge around 50 episodes but could take as many as 200 episodes. For $N=3$ we have only seen it converge once within 500 episodes, as shown in Fig. 2. So the actual performance of these algorithms with respect to convergence seems to depend more on neural network initialization or the randomly generated dataset, which is not a great property for an algorithm.
@@ IV-D Discussion
Figure 6: RMSE analysis for a simulated platoon of 100 vehicles (blue: DMPC controller, orange: neural network controller, green: linear feedback controller). The zoomed in portion of the graph enables us to see the tradeoff between the feedback controllers and a DMPC controller.
One potentially useful application of our algorithm is the ability to use the resulting controllers to guide control design. Figure 7 shows the level sets of the learned controller we used in simulation and the level sets of the linear feedback controller used. The level sets of the neural network controller paint a clear picture about what the neural network learned to value.
When a vehicle is too close to and is traveling faster than its predecessor (i.e., both position error and velocity error are less than 0, the bottom left quadrant), the controller’s outputs quickly ramp up to return the maximum deceleration of -3. Similarly, the controller quickly corrects based on velocity error in the top left and top right quadrants. On the other hand, when a vehicle is too far away but traveling faster than the predecessor (the bottom right quadrant), there is less urgency. This is because this scenario is both conducive to stability (position error will decrease) and is the most safe. Thus, we see that unless the vehicle’s state is in the bottom right quadrant, the controller heavily weighs the velocity error as opposed to the position error.
Considering these qualitative observations, the level sets appear to approximate a piecewise-linear function before saturating at the control limit. Future research could try to use this information to learn saturating piecewise-linear controllers that are provably stable via stability analysis similar to [9] instead of via a learned Lyapunov certificate.
Figure 7: Level sets for the learned neural network controller and the linear feedback controller studied. Clearly, the learned controller is a nonlinear function of the two-dimensional input.
V CONCLUSIONS
@@
We present an algorithm to learn a controller for dynamic systems. In cases where the state dimension is small, the algorithm permits learning a Lyapunov function that guarantees stability of the closed-loop system. The benefit of our algorithm is the option to design an loss function that guides the controller to behave in a desired way. We apply the algorithm to a platooning problem with an intrinsic tradeoff between traffic throughput/fuel efficiency and vehicle safety. Due to a change of variable, we can learn a single controller for a platoon of heterogeneous vehicles, and we investigate the performance of a controller learned using our algorithm on a platoon of four F1Tenth vehicles and in a simulated environment with 100 vehicles. We show that we can train an algorithm in simulation and obtain performance that bridges the gap between DMPC (the state-of-the-art when vehicles can communicate with one another) and classical linear feedback controllers.
@@
Future work will consider explicit safety certificates (barrier functions) along with the stability certificates. Further, it is possible that encoding some different notion of state, either by using a history of past errors as input into the neural network or using a recurrent neural network, could lead to better performance. We hope to investigate this using either this algorithm or similar (e.g., reinforcement learning) to investigate the performance limits of feedback controllers for platoons using a CDH spacing policy and a PF topology.
ACKNOWLEDGMENT
@@
Research was sponsored by the DEVCOM Analysis Center and was accomplished under Cooperative Agreement Number W911NF-22-2-001. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.

### v2

@@ 4.1 Controller learning
4.2 Baseline controllers
For the hardware and simulation experiments, we compare against a linear feedback and a DMPC controller. The linear feedback controller computes control inputs using
$u_{i}^{\text{lin}}(t)=k^{T}x_{i}(t)$
where $k\in\mathop{\mathbb{R}}^{2}$ is a vector of gains and $x_{i}(t)$ is vehicle $i(t)$’s error state at time $t$. We used $k=(1,2)$ for all experiments as we have previously found these gains work well on the F1Tenth vehicles [23]. The DMPC controller assumes each vehicle receives the planned state trajectory of the vehicle in front of it. Then at each timestep $t$, each vehicle solves an optimization problem that minimizes a weighted objective that penalizes deviation from its predecessor’s plan, deviation from its own plan calculated at the previous timestep, and the control input size. See [23, 22] for implementation details and further analysis of the DMPC controller.
4.3 Hardware experiments
To validate our algorithm on hardware, we use the F1Tenth platform, modified as described in [23]. We test the vehicles in a roughly 4 m by 8 m oval racetrack, shown in Fig. 1. We model each vehicle’s dynamics using Eq. 1 with $\tau_{i}=0.3$ (found using a least-squares regression on experimental data [16]). In all experiments, the lead vehicle uses the quadratic cost function DMPC controller described in [23] to track a safe velocity based on the current curvature of the course. This leads to faster trajectories during the straight portions of the course and slower trajectories around the curves. Each following vehicle uses the neural network controller learned using algorithm 1 with a desired distance of 0.75 m.
To ensure safety of the learned controller, we use $L_{\text{safe}}$ to penalize when a vehicle has a distance of less than 0.25 m to its predecessor. To encourage smooth trajectories with $L_{\text{slew}}$, we penalize large actions and large changes in speed from one timestep to the next. We also add two ReLU layers and a linear layer to the neural network to ensure the controller only outputs commands such that $\left|u_{i}\right|\leq 3\,\text{ m/s}^{2}$. To guide the controller toward closed-loop stability, we use ${\cal L}_{\text{stab}}$ to penalize a running sum of the difference in error from the next timestep to the current timestep over the horizon (because we ideally want the errors to decrease). Other choices for $L_{\text{stab}}$ can try to guide the controller towards improving the convergence rate of the system to the origin. We ran algorithm 1 with $N=5$, where both $V$ and $\pi$ are two-layer feedforward neural networks with 8 neurons in each hidden layer and leaky ReLU activation functions with slope 0.1 for around 100 episodes to learn a controller.
@@ 4.3 Hardware experiments
Based on the results in Figs. 3 and 4, it is difficult to determine which algorithm performs the best. We do note that when using the neural network controller, zero collisions occurred between vehicles or with the course boundaries. With the other two controllers, however, vehicles would sometimes collide with the course boundaries. Additionally, with the linear feedback controller, vehicles would sometimes collide with one another when braking at the end of the experiment. We speculate that the collisions with the boundaries occurred because the controllers attempted to follow the predecessor’s velocity profile too aggressively when rounding the curves, leading to lateral slipping.
4.4 Simulation experiments
@@
For a platoon using a PF topology and a CDH spacing policy, ensuring safety becomes more difficult as the platoon size increases [21, 28, 12]. To investigate the limits of our approach, we simulate a platoon of $N=100$ vehicles and train a controller using a desired distance of 5 m. We use an almost identical setup as in the hardware section, except we use two hidden layers with 32 neurons each, we penalize a vehicle for getting within 2 m of its predecessor, and we use a heterogeneous platoon where the dynamics parameter $\tau_{i}\in[0.2,0.8]$ is selected at random. As mentioned in the previous section, we cannot verify the Lyapunov certificates because 1) the search space becomes very large and 2) the number of binary variables required to solve Eq. 7b increases linearly with the platoon size leading to potentially exponential increase in the time required to solve the MILPs. However, we still find that training for a small amount of time (around 100 simulated episodes) yields controllers that perform well in test scenarios.
@@
As shown in [23], there is a large gap in the performance of DMPC controllers and linear feedback controllers as platoon size scales. Though DMPC should outperform feedback methods due to the ability to communicate planned trajectories, it would be useful in practice if a feedback controller can perform similarly well. For example, in the scenario where communication is temporarily lost, it would be beneficial to have a feedback controller that can take the place of the DMPC method without disrupting the platoon’s motion. Thus, we again compare our learned controller to DMPC and linear feedback, with the goal of bridging the gap between DMPC methods and linear feedback methods.
@@
Figure 5: Simulated trajectories for a platoon of 100 vehicles when using the learned neural network controller. The lead vehicle tracks the velocity profile shown by the reference. The faded lines in the position plot depict the desired position of each vehicle.
For the simulation experiments, the first vehicle tracks a velocity profile that starts at 20 m/s, accelerates to 25 m/s, then decelerates back to 20 m/s. We add Gaussian noise to both the forward dynamics and to the sensing of the position/velocity errors to ensure the learned controller still works in a noisy setting. Note that for a vehicle traveling 20 m/s, common practice would suggest a spacing of 40 m [27], so a 5 m spacing is aggressive for this scenario. The simulation results for the neural network controller are shown in Fig. 5. As expected when using a feedback controller with a PF topology, the errors do propagate down the platoon. However, the controller does stabilize the platoon to the true desired velocities quickly and without any collisions within the platoon.
@@
Figure 6: RMSE analysis for a simulated platoon of 100 vehicles (blue: DMPC controller, orange: neural network controller, green: linear feedback controller). The zoomed in portion of the graph shows the differences between the feedback and DMPC controllers.
4.5 Discussion
The algorithm proposed in this work to concurrently learn a neural network controller and a neural network Lyapunov function achieves its goal of generating a controller that significantly outperforms hand-designed controllers. Furthermore, the learned controller translates well from simulation to the real world. However, there is still a need to improve this algorithm (and others proposed in prior works) to achieve convergence for high-dimensional systems, i.e., beyond a state dimension of around ten. From a practical perspective, however, the learned controller typically performs well after only a few tens of episodes, even for very large platoons.
To try to improve the convergence rate of our algorithm, we attempted to “warm-start” the Lyapunov function by sharing weights from a previously verified Lyapunov function with fewer vehicles $N$. However, we did not find that this improved performance. Instead, convergence rates were inconsistent even for small $N$. For $N=2$, the algorithm would typically converge within 50 episodes but could take as many as 200 episodes. For $N=3$, the algorithm would typically converge within 500 episodes, but could take as many as 1000 episodes. Thus, the actual performance of these algorithms with respect to convergence seems to depend on neural network initialization or the randomly generated dataset. Future research will investigate how to improve the convergence rate.
5 Conclusion
@@
We present an algorithm to learn a controller for dynamic systems. When the state dimension is small, the algorithm returns a Lyapunov function that guarantees stability of the closed-loop system. The benefit of our algorithm is the option to design a loss function that guides the controller to behave in a desired way. We apply the algorithm to a platooning problem with an intrinsic tradeoff between traffic performance and safety. Due to a change of variable, we can learn a single controller for a platoon of heterogeneous vehicles, and we investigate the performance of a controller learned using our algorithm on hardware with a platoon of four F1Tenth vehicles and in a simulated environment with 100 vehicles. We show that we can train an algorithm in simulation and obtain performance that bridges the gap between DMPC (the state-of-the-art when vehicles can communicate with one another) and classical linear feedback controllers.
@@
Future work will consider explicit safety certificates (barrier functions) along with the stability certificates. Furthermore, it is possible that encoding some different notion of state, either by using a history of past errors as input into the neural network or using a recurrent neural network, could lead to better performance. We hope to investigate this using either this algorithm or similar (e.g., reinforcement learning) to investigate the performance limits of feedback controllers for platoons using a CDH spacing policy and a PF topology.
Acknowledgements
@@
Research was sponsored by the DEVCOM Analysis Center and was accomplished under Cooperative Agreement Number W911NF-22-2-0001. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the DEVCOM Analysis Center or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.